OpenStudy (anonymous):

x · 1 + = _____. 1 x 1 + x 2x

OpenStudy (anonymous):

Hold on

OpenStudy (ahsome):

\(x*1=x\) Kinda like how \(10*1=10\)

OpenStudy (ahsome):

All numbers, times by 1, will equal that same number

OpenStudy (ahsome):

All I see is �����������

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

xX1+x/1=

OpenStudy (ahsome):

So, equatiom: \[x*1+\frac{x}{1}\]?

OpenStudy (anonymous):

yes

OpenStudy (ahsome):

\(x*1=x\) Therefore, we can simplfy: \[x*1+\frac{x}{1}\]\[x+\frac{x}{1}\] All numbers, that are not a fraction, can have a one as a denominator \[\frac{x}{1}+\frac{x}{1}\] Add the two numbers together \[\frac{2x}{1}\] Any number, divided by 1, equals the same number. Therefore, the answer is: \[2x\]

OpenStudy (ahsome):

That helpful?

OpenStudy (anonymous):

Thank you for explaining and not just giving the answer